OpenStudy (anonymous):

Find the vertices and foci of (x+2)over 4^2+(y-1)0ver 25^2=1

OpenStudy (anonymous):

\[(x+2)^{2}/4+(y-1)^{2}/25=1\]

OpenStudy (anonymous):

@ShailKumar

OpenStudy (anonymous):

This is an ellipse and the vertex is (-2,1)

OpenStudy (anonymous):

for the foci i got (-2, 1+or- square root 21)

OpenStudy (anonymous):

and for the vertices i got (-2,6) (-2,-4) but when i put it in a graphing calculator i get different vertices and foci

OpenStudy (anonymous):

https://www.desmos.com/calculator/46lmaqxorx

OpenStudy (anonymous):

(2,-1) is the center of the ellipse. How did you find the vertices ?

OpenStudy (anonymous):

I used the formula (h, K+a) (h,k-a) because the major axis is vertical

OpenStudy (anonymous):

(-2,1) is the center of the ellipse

OpenStudy (anonymous):

Actually you need to calculate eccentricity first

OpenStudy (anonymous):

its parallel to the y-axis because a is always biggest number and its under y^2

OpenStudy (anonymous):

oh, is that where e=c/a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is c square root 21 though?

OpenStudy (anonymous):

because then e=square root 21/5

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

ok why is eccentricity important in finding the vertices and foci?

OpenStudy (anonymous):

If center is (h,k), then foci are (h, k+c) and (h, k-c)

OpenStudy (anonymous):

ok thats what i did and i got (-2,1+square root 21) (-2,1-square root 21)

OpenStudy (anonymous):

eccentricity decide which conic section (parabola, ellipse or hyperbola) you are talking about. So their properties should be dependent on eccentricity.

OpenStudy (anonymous):

Nice Question ShailKumar

OpenStudy (anonymous):

You did it right for foci.

OpenStudy (anonymous):

ok maybe i was interpreting the graph incorrectly

OpenStudy (anonymous):

BTW, for this ellipse e = c/b. That plot is showing a parabola.

OpenStudy (anonymous):

on the graph the vertices were like (-4,0) and (0,0) though

OpenStudy (anonymous):

isnt e=c/a not b?

OpenStudy (anonymous):

Forget that. you've done correctly.

OpenStudy (anonymous):

oooh the graph on the link was showing a parabola?! ok that explains a lot

OpenStudy (anonymous):

Could you find vertices?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I thought they were (-2,6) and (-2,-4)

OpenStudy (anonymous):

but then those points arent on the x-axis...

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

Vertices are (h,k+b) and (h,k-b).

OpenStudy (anonymous):

arent they (h,k+a) (h,k-a)?

OpenStudy (anonymous):

In your equation of ellipse: you have taken a^2 = 25 and b^2 = 4. right ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Then you are right. It should be (h,k+a) (h,k-a).

OpenStudy (anonymous):

BTW, you've got correct vertices (-2,6) and (-2,-4)

OpenStudy (anonymous):

it looks so weird lmost like a circle when i graph it though

OpenStudy (anonymous):

and it doesnt make sense because then all of the points are on x=-2 and how am i supposed to draw that? i dont know any x-coordinates

OpenStudy (anonymous):

https://www.desmos.com/calculator/fwn45vpf2z this graph is showing different vertices

OpenStudy (anonymous):

ok yea that one was a parabola graph i just posted a new one though and it still shows different vertices

OpenStudy (anonymous):

You have interchanged the roles of a and b.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

In the plot you've got, you have put a>b. But in the problem a<b

OpenStudy (anonymous):

a=25 and b=4 so a>b always in an ellipse

OpenStudy (anonymous):

Not always. You post standard ellipse equation.

OpenStudy (anonymous):

oh ok so im just confused about how to graph this problem because how do i know how wide the ellipse is when i dont have any x-coordinates?

OpenStudy (anonymous):

For width you don't need x coordinates. Actually width of this ellipse is 2a = 4. Length is 2b = 10 !

OpenStudy (anonymous):

Did you see the attached image?

OpenStudy (anonymous):

which attached image?

OpenStudy (anonymous):

Attaching again

OpenStudy (anonymous):

Ummm I have (x+2)^2/4+(y-a)^2/25=1 is this the same as you have?

OpenStudy (anonymous):

Sorry thats (y-1)^2/25

OpenStudy (anonymous):

Actually I have (x+4)^2/4+(y-1)^2/25=1.

OpenStudy (anonymous):

it should be (x+2)^2/4

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

im still confused about how to graph it...I can plot the vertex, the vertices but how can i draw an ellipse without knowing how wide to draw it?

OpenStudy (anonymous):

Do you wanna draw it by hand ?

OpenStudy (anonymous):

yes i have to draw it by graphing paper

OpenStudy (anonymous):

how do i plot 2a=4 and 2b=10?

OpenStudy (anonymous):

Ellipse has two perpendicular axes called major axis and minor axis. First, check a>b or b>a. If a>b, major axis is parallel x axis otherwise parallel to y axis.